# LLaVA-3D: A Simple yet Effective Pathway to Empowering LMMs with 3D-awareness

### v1

@@ 5.7 More Analysis
To better understand the impact of different components of our LLaVA-3D, we conduct a thorough ablation study on the ScanQA and SQA3D benchmarks.
Impact of Pooling Strategy.
Given the maximum token length limitation of LLMs, we apply pooling to the 3D patch tokens extracted from multi-view images to reduce the number of tokens. However, this pooling process inevitably leads to some information loss. To understand its impact on performance, we conducted experiments to evaluate the effects of pooling. As shown in Tab. 10, the voxelization pooling strategy outperforms the FPS pooling method on 3D QA benchmarks. Reducing the voxel size in the voxelization pooling or increasing the number of 3D patch tokens obtained through FPS pooling can both enhance the model’s performance to some extent.
@@
Pooling Stratgey Voxel Size Token Number ScanQA SQA3D
@@
Multi-View Images Sampling Strategy.
Due to the redundancy of information among multi-view images and considering computational costs, we sample $V$ views from the egocentric images of the 3D scene in our experiment. In this section, we explore the effect of different image sampling strategies during the inference stage: 1) Uniform Sampling: To achieve comprehensive coverage of the entire 3D scene, a straightforward approach is uniform sampling, which samples images evenly. 2) Text-Guided Sampling: such sampling strategy uses CLIP to select the frames related to the text instruction during inference based on the image-text CLIP similarity score. Our experiments demonstrate that these two sampling strategies share similar performance on ScanQA and SQA3D, so we choose uniform sampling for simplicity.
Number of Views.
An intuitive assumption is that sampling more views from the 3D scene will preserve more information about the 3D scene. We conduct a comparative experiment varying the number of views sampled from 3D scenes. Tab. 11 presents the Exact Match (EM) scores on ScanQA and SQA3D across different settings, revealing that the increase in EM score is marginal as the number of views increases. Additionally, the experimental results indicate that exceeding a certain number of views can degrade the model’s performance.
@@
6 Conclusion
We propose LLaVA-3D, a simple yet effective framework built upon the well-established 2D LLaVA model. LLaVA-3D extends the LLaVA’s capabilities to understand the 3D world by using 3D patches to bridge 2D features within a 3D space, enabling spatial understanding while efficiently preserving the original 2D image understanding and reasoning capability. Experimental results show that our method sets new state-of-the-art performance on a wide range of 3D tasks and benchmarks. We hope that our model will inspire new ideas for building 3D LMMs, and in the future, we plan to explore the application of LLaVA-3D in more downstream scenarios, such as robot manipulation and navigation.

### v3

@@ 5.5 More Analysis
In this section, we first delve deeper into the architectural benefits and efficacy of adapting the 2D LMM to 3D, as opposed to developing a 3D LMM solely from LLMs. Then, we analyze the effectiveness of the components and validate the generalizability of our method on various 2D LMMs.
Developing 3D LMM from LLM or 2D LMM. As shown in Tab. 8, (a), (b), and (c) develop 3D LMM from the LLM with different 3D scene feature and 3D-language connectors, while (d) builds 3D LMM from well-trained 2D LMM: LLaVA-1.5 [34]. Due to the context length limitation of the LLM in (b), (c), and (d), we utilize 3D voxelization pooling to compress the 3D patch tokens. The results demonstrate: 1) the Q-Former (a) and Pooling + MLP (b) share a similar performance on 3D V&L benchmarks. 2) using CLIP (c) alone instead of SAM + CLIP (b) like 3D-LLM [15] achieves better performance and significantly reduces 3D scene feature construction time from 900s to 0.2s. 3) Developing from 2D LMM (d) instead of LLM (c) could greatly improve the 3D scene understanding performance, achieving improvements of 4.1% Acc@0.25 on ScanRefer.
Effectiveness of 3D Patch. To further ascertain the effectiveness of our proposed 3D Patch, we conduct additional experiments across a variety of 3D question answering and 3D dense captioning benchmarks. As shown in Tab. 9, injecting 3D position information into 2D patches only brings minor improvements on ScanQA and SQA3D benchmarks. However, 3D patches prove instrumental in tasks requiring objects position information and more difficult 3D spatial understanding, yielding substantial performance gains of 13.3% on the MMScan QA benchmark and a remarkable 54.4% improvement on the Scan2Cap benchmark.
Effectiveness of Grounding Decoder. An ideal scenario for the 3D LMM to perform 3D visual grounding would be to directly output 3D bounding boxes in text format, just like the 2D LMM. However, while the 3D LMM can accurately output the object dimensions represented by a given 3D coordinate, it struggles to directly output the 3D position of the target object. We attempt to train the LLaVA-3D model to directly output the coordinates of the target object in the 3D visual grounding task in text format or using special tokens [15], resulting in 7.8 Acc@0.25 and 8.2 Acc@0.25 on the ScanRefer benchmark. However, utilizing our grounding decoder could achieve a performance of 50.1 Acc@0.25.
Generalization to various 2D LMMs. In Tab. 8, we adapt our LLaVA-3D to other 2D LMMs, such as LLaVA-1.5 [34] (d) and InternVL2.5 [12] (e). Experiments demonstrate that our method could be a general extension to equip 2D LMMs with 3D capabilities due to the simple yet effective designs. Comparing (d) and (e), we can observe that LLaVA-3D can benefit from stronger 2D understanding capabilities of the 2D LMM base model. Besides, due to the intrinsic consistency between multi-view image 3D scene representation and video, the results in (e, f) demonstrate that our method can enjoy the significant benefits from video LMMs.
6 Conclusion
We propose LLaVA-3D, a simple yet effective framework built upon the well-established LLaVA. LLaVA-3D extends LLaVA’s capabilities to perform 3D scene understanding and grounding by leveraging the 3D patches and grounding decoder while efficiently preserving the 2D visual understanding and reasoning capability. Experimental results show that our method achieves state-of-the-art performance on various 3D tasks and benchmarks. We hope that our model will inspire new ideas for building 3D LMMs, and we plan to explore the application of LLaVA-3D in more downstream scenarios, such as robot manipulation and navigation.
Supplementary Material
Appendix A Implementation Details
LLaVA-3D is built upon the LLaVA-Video-7B [53], utilizing their pre-trained weights from the HuggingFace library, and follows a two-stage training process. Each subsequent stage builds upon the weights learned in the previous stage. The number of views $V$ is set to 32. When adapting our method to LLaVA-1.5 [34], due to the LLM context length limitation, we use the voxelization pooling to compress the 3D patch token numbers, and the maximum number of 3D patch tokens after 3D pooling is set to 3096. All experiments are conducted on 16 × 80G A100 GPUs.
Settings of Stage 1. We use the Adam optimizer to train our model for one epoch with a total batch size of 16 and a warmup ratio of 0.03. During the warmup phase, the learning rates peak at 1e-5 for the LLM, 3D position encoding layer and grounding decoder, and 2e-6 for the vision encoder. The training objectives consist of the auto-regressive language modeling loss and the grounding decoder training loss.
Settings of Stage 2. In stage 2, we freeze all the components except for the grounding decoder. The model undergoes 40 training epochs on 16 A100 GPUs with a peak learning rate of 1e-4.
Appendix B Training Convergence Speed
To further validate the effectiveness of 2D LMM-based Architecture and ensure fairness as much as possible, we choose LLaVA-1.5 as the base model and replace the LLaVA-3D-Instruct-86K dataset in stage 1 with the MMScan QA [38] training data. We record and evaluate the performance of LLaVA-3D under different training data ratios. Besides, we further fine-tune LEO [18] on full MMScan QA training data based on the officially released model checkpoint. Both models utilize Vicuna-7B as the LLM, ensuring comparable parameter counts. As illustrated in Fig. 5, LLaVA-3D surpasses LEO’s full-step performance even when trained on less than 300 steps, indicating better data efficiency and 3.5x faster training convergence speed.
Figure 5: Training convergence comparison. LLaVA-3D achieves higher data efficiency and faster convergence speed during the instruction tuning stage compared with existing 3D LMM: LEO.
Appendix C More Architecture Details
In this section, we provide more details about how to connect LLM with the grounding decoder via special token embeddings, which enables the end-to-end optimization of the entire model and the grounding decoder details.
C.1 Connecting with Grounding Decoder
For 3D understanding tasks like 3D visual grounding, we employ the grounding decoder to localize objects according to the user query. Specifically, we introduce a special localization token <LOC> into the LLM vocabulary, and the LLM is trained to predict the special location token to represent the 3D bounding boxes prediction when the task necessitates 3D bounding box outputs. The last layer embedding of this location token is then sent to the 3D grounding decoder as a condition. Grounding Decoder receives both the 3D patch features and the obtained location token embeddings as inputs and predicts the 3D visual grounding results.
C.2 Grounding Decoder Details
Here illustrate more architectural and training objective details about the grounding decoder. Our grounding decoder consists of $L=4$ decoder layers, as illustrated in Fig. 6. For query initialization, we employ farthest point sampling to select $N=512$ instance queries from the 3D patches. We initialize the value of queries to zeros and only the 3D coordinates of the sampled queries are used to set the corresponding learnable positional encoding.
Multi-Scale 3D k-NN Cross Attention.
During the cross-attention between the queries and 3D patch features, the instance queries can only attend to the features of $k$ nearest 3D patch neighbors to accelerate training convergence speed and reduce the memory usage. To capture the object geometry information at different scales, the $k$ is set to be $\{16,32,64,128\}$ in different decoder layers. Inspired by ODIN [21], we encode the relative position between the query and its neighbor into the position encoding with a learnable MLP. The position encoding is incorporated into the cross-attention computation by adding it to the instance queries and their corresponding 3D patch neighbors.
Distance-Adaptive Self Attention.
After attending to the 3D patch features, we utilize the distance-adaptive self-attention layer [13] to model the relative spatial relationship among the queries and achieve the visual-language feature communication. Distance-adaptive self-attention introduces a bias term based on the distances between instance queries. The pairwise attention between the i-th instance query and the j-th instance query is computed as:
$Attn(Q_{i},K_{j},V_{j})=Softmax(\frac{Q_{i}K_{j}^{T}}{\sqrt{C}}-\sigma\cdot D)V_{j},$ (2)
where $Q,K,V$ is the query, key, and value of the attention module, $C$ is the channel of the embedding, $\sigma$ is a learnable parameter based on the query, and $D$ indicates the Euler distance between the position of these two instance queries. For the location token that does not have the 3D coordinate information, the bias term between this location token and the instance queries is set to zero.
Box Head.
The box head consists of a two-layer MLP, which takes the updated instance queries in each decoder layer as input and predicts the corresponding 3D bounding boxes.
C.3 Training Objective
After matching the instance queries with ground truth 3D bounding boxes, for each match between a proposal and a ground truth object, we compute the DIOU loss [11] between predicted and ground truth boxes. We utilize InfoNCE loss [41] to optimize the similarity between the matched queries and the location token.
Figure 6: Grounding Decoder Architecture.
Appendix D More Components Analysis
@@
To better understand the impact of different components and the generalizability of our LLaVA-3D, we conduct a thorough ablation study on the ScanQA and SQA3D benchmarks based on LLaVA-1.5 [34].
Impact of Pooling Strategy. Here we conduct various experiments to evaluate the effects of the different pooling strategies. For voxelization pooling, we adopt the simple voxelization approach from ODIN [21]. As shown in Tab. 10, the voxelization pooling strategy outperforms the FPS pooling method on 3D QA benchmarks. Model performance can be improved by either decreasing voxel size in voxelization pooling or increasing the number of 3D patch tokens in FPS pooling.
@@
Pooling Strategy Voxel Size Token Number ScanQA SQA3D
@@
Multi-View Images Sampling Strategy. To balance computational efficiency with visual coverage, we sample $V$ views from the egocentric images of each 3D scene. We investigate two sampling strategies during inference: Uniform Sampling, which evenly samples images across the scene, and Text-Guided Sampling, which selects frames based on CLIP image-text similarity scores to the input instruction. Since our experiments show a similar performance, we adopt uniform sampling for its simplicity.
Number of Views. An intuitive assumption is that sampling more views from the 3D scene will preserve more information about the 3D scene. We conduct a comparative experiment varying the number of views sampled from 3D scenes. Tab. 11 presents the Exact Match (EM) scores on ScanQA and SQA3D across different settings, revealing that the increase in EM score is marginal as the number of views increases. Additionally, the experimental results indicate that exceeding a certain number of views can degrade the model’s performance.
@@
Appendix E More Qualitative Results
3D Scene Understanding.
We evaluate LLaVA-3D on various 3D scene understanding tasks and display more visualization results from Fig. 7 to Fig. 9. These examples demonstrate LLaVA-3D’s robust 3D understanding abilities: comprehensive 3D scene understanding, accurate object recognition, and precise object localization in the 3D world. Besides, our model enables the users to more easily interact with the 3D scene through the 2D images.
Appendix F Video Demo Comparision
To enhance real-world applicability, we design our framework to process 2D videos - a widely accessible data format that users can capture with standard mobile devices. Our pipeline processes these inputs by uniformly sampling 32 frames and leveraging DUST3R [48], an efficient offline MVS method, to obtain depth maps, camera parameters, and poses. Notably, DUST3R completes this process within one minute, enabling seamless conversion of conventional video inputs into our model’s required format. To validate our approach, we conduct comprehensive evaluations against LLaVA-OneVision 72B [26] in Fig. 10, a state-of-the-art multimodal model that demonstrates strong capabilities across diverse 2D scenarios, including single-image understanding, multi-view reasoning, and video understanding. The qualitative results reveal that our method achieves superior performance in 3D spatial reasoning and relationship understanding between objects with significantly fewer parameters (7B), highlighting the effectiveness of our 3D-aware architecture.
Figure 7: LLaVA-3D could perform 2D Click-based 3D dense captioning, generating the corresponding object caption and 3D bounding box.
Figure 8: LLaVA-3D could perform 2D Click-based 3D question answering, now users could click on the 2D images and ask the question.
Figure 9: LLaVA-3D exhibits powerful 3D visual grounding capability, enabling accurate 3D bounding boxes output.
Figure 10: LLaVA-3D achieves superior performance in 3D spatial reasoning and relationship understanding between objects with significantly fewer parameters compared with powerful LLaVA-OneVision 72B.